One number is two less than five times another. If their sum is increased by two, the result is twenty-four. Find the numbers.

Answer by JulietG(1812)

A = 5B-2 (One number is two less than five times another.)
A + B = 22 (If their sum is increased by two, the result is twenty-four.)
Take the known value of A from the first equation and put it into the second.
A + B = 22
(5B-2) + B = 22
6B - 2 = 22
Add 2 to each side
6B = 24
Divide each side by 6
B = 4
If the smaller number is 4, then the larger number is (5*4)-2, or 18
18 + 4 = 22